OpenStudy (anonymous):

f(x)=x4-2x find f(2i) how do i solve this

jimthompson5910 (jim_thompson5910):

f(x) = x^4 - 2x f(2i) = (2i)^4 - 2(2i) f(2i) = 2^4*i^4 - 2(2i) f(2i) = 16(1) - 2(2i) f(2i) = 16 - 4i

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

8/1-2i